UsernameTaken:

Jonathan sets a camera on a tripod to film birds in a nest. The camera is 1.8 m off the ground and 16.4 m away from the tree. The angle of elevation from the camera to the bird’s nest is 29°. How far above the ground is the bird’s nest? (I accidentally closed this one last time)

chuckthedonut:

Are there any options?

UsernameTaken:

Yep, the options are 7.3, 9.1, 10.9, and 14.3. I thought I had it at first but I'm still confused.

Narad:

The distance \[= 1.8 + 16.4 \tan29º\]

NotLeiland:

You need to use S\[\frac{ O }{ H }\] (Opposite over Hypotenuse) C\[\frac{ A }{ H }\] (Adjacent over Hypotenuse) T\[\frac{ O }{ A }\] (Opposite over Adjacent) Since this question gives you the Opposite and the Adjacent angles, all you need to do is do 1.8 + 16.4tan29 in your calculator and you should have your answer.